OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x – y = 10

OpenStudy (anonymous):

So the equation is x+3y=6, and you want to solve for y? Just move everything that is not a y to the other side of the equals sign. Standard procedure.

OpenStudy (anonymous):

y = (6 - x)/3

OpenStudy (anonymous):

@xxmileyxxd This "so y=6-x/3, or y=-x+6/3 so your left with y=2-x or y=-x+2" Doesn't make sense. Can you explain further, please?

OpenStudy (anonymous):

xxmileyxxd is incorrect. You would not get "y=2-x or y=-x+2" remember it's (6 - x) divided by 3. you cannot divide the 6 alone by 3. you need to divide the "x" as well. so it would be: \[y = 2 - \frac{ 1 }{ 3 }x \]

OpenStudy (anonymous):

He's just solving for y, you move everything to one side to get your answer, than divide the 3 to the other side.

OpenStudy (anonymous):

^ yes, which would yield the following: \(\large x + 3y = 6 \rightarrow y=\frac{6-x}{3}\) In slope intercept form that would be \(\large y=-\frac{1}{3}x+2\).